Determine if the two equations are parellel, perpendicular or neither
2/5x - 1/3y = 7/8
1/3x - 2/4y = -72
This is not from a book

Answer by stanbon(48554)

Determine if the two equations are parellel, perpendicular or neither
2/5x - 1/3y = 7/8
1/3x - 2/4y = -72
---------
Put each equation in slope-intercept form:
y = ((2/5)/(1/3))x -[(7/8)/(1/3)]
y = (6/5)x -(21/8)
------------
y= [(1/3)/(1/2)]x + [72/(1/2)
y= (2/3)x +144
-------------
They are neither parallel or perpendicular as they have different slopes
and different y-intercepts.
-------------.
Cheers,
Stan H.